### v1

# A QUBO Formulation for the Generalized LinkedIn Queens Game

Abstract
In this paper, we present a QUBO formulation designed to solve a series of generalisations of the LinkedIn queens game, a version of the N-queens problem. We adapt this formulation for several particular cases of the problem by trying to optimise the number of variables and interactions, improving the possibility of applying it on quantum hardware by means of Quantum Annealing or the Quantum Approximated Optimization Algorithm (QAOA). We also present two new types of problems, the Coloured Chess Piece Problem and the Max Chess Pieces Problem, with their corresponding QUBO formulations.
1 Introduction
The N-queens problem [1] is a problem that has attracted enormous attention since its proposal in 1848 by the chess player Max Bezzel, leading many mathematicians, including Gauss, to try to solve and generalise it. This problem can be solved efficiently, while the problem of determining how many solutions exist becomes much more complicated. Such a problem has no applications, but it is a good problem on which to test methods of solving logically constrained problems.
LinkedIn recently published its Queens game (which we will call LQueens), based on N-queens problem, but with some changes in the restrictions to be carried out and the initial conditions. In this new version, the restriction on diagonals is eased, but a restriction on regions is imposed.
In recent years, quantum computing has gained great popularity due to its different computational capabilities compared to classical computing, making it possible to tackle certain problems more efficiently. Some of these problems are combinatorial optimisation problems, in which quantum annealing using quantum annealers [2, 3] such as those of Dwave and the Quantum Approximated Optimization Algorithm (QAOA) [4] using digital quantum computers stand out. In both cases, the formulation of the problem comes through a quadratic unconstrained optimisation problem (QUBO) [5], which contains the interactions to be optimised between the different variables, represented by logical qubits, minimising the energy of the system to find the lowest cost combination. These methods can be adapted to cases with constraints by adding terms that increase the energy substantially if the constraints are not respected.
Recently, the N-queens problem has been solved with a quantum algorithm in a quantum simulator [6], but not generalized and formulated in an optimal formulation for the currently most standardized algorithms. On the other hand, there are also QUBO formulations for other games such as Sudoku [7].
Motivated by this new version of the game and the quantum computing algorithms, we present a series of generalisations of the N-queens problem and a QUBO formulation associated with each of them so that they can be solved by quantum means, such as quantum annealing or the Quantum Approximated Optimization Algorithm (QAOA). With these generalizations, we can define two new types of problems, which take advantage of the definitions and terms we created for the N-queens generalizations.
2 LinkedIn’s Queens and N-Queens problem
The original problem to be solved in the N-queens consists of placing $N$ queens on a chessboard of dimensions $N\times N$ in such a way that none of them threatens the others. That is, no queen can be in the same row or column as another queen, nor be diagonal to another queen. The constraints are shown in Fig. 1.
Figure 1: Constraints on a $5\times 5$ problem, such that no queens can be placed in any of the cells marked in red. a) Column constraint, b) Row constraint, c) Diagonal constraint on N-queens, d) Diagonal constraint on LQueens, e) Region constraint.
To formalize it mathematically, we say that the state of the board is determined by a binary matrix $x$ whose elements $x_{ij}$ determine the state of the cell in row $i$ and column $j$, so that if it is 0 there is no queen in that cell, and if it is 1 there is. Therefore, the restrictions are
  • Rows condition: there can only be one queen per row (Fig. 1 a):
    $\sum_{j=1}^{N}x_{ij}=1\quad\forall i\in[1,N].$
  • Columns condition: there can only be one queen per column (Fig. 1 b):
    $\sum_{i=1}^{N}x_{ij}=1\quad\forall j\in[1,N].$
  • Diagonal condition: there cannot be two queens in the same diagonal (Fig. 1 c):
    $\sum_{(k,l)\in D_{i,j}}x_{kl}\leq 1\quad\forall i,j\in[1,N],$
    being $D_{ij}$ the set of cells on the diagonals of cell $(i,j)$, including it.
The condition that there be $N$ queens is implicit in the first two restrictions.
LinkedIn’s version of the game is slightly different. While it retains the first two restrictions, the third restriction becomes softer, making it so that queens can now not be diagonal only in the case where they are adjacent, which translates to
  • Diagonal condition: there cannot be more than one queen in the same adjacent diagonal (Fig. 1 d): $x_{ij}+X_{i+1,j+1}+X_{i-1,j+1}+X_{i+1,j-1}+X_{i-1,j-1}\leq 1\quad\forall i,j\in[1,N]$, being $X_{kl}=x_{kl}$ if $k,l\in[1,N]$ and 00 otherwise.
Another modification in the LinkedIn version is the region restriction, which imposes on us that we have to place one and only one queen for each one of the $N_{R}=N$ colored regions, which translates to
  • Region condition: there can only be one queen per region (Fig. 1 e): $\sum_{(i,j)\in R_{k}}x_{ij}=1\quad\forall k\in[1,N_{R}]$, being $R_{k}$ the cells belonging to the $k$-th region.
An additional condition in this case is that there are a number $N_{I}$ of queens already placed in cells $A_{I}$, as an initial condition. This can be addressed by forcing the variables associated with these cells to be forcibly 1, which leads to all other adjacent variables in the same row, in the same column and in the same region to be forcibly 0. This translates to
  • Initial condition: there are $N_{I}$ pre-placed queens:
    • Cell initialized: $x_{lm}=1\quad\forall(l,m)\in A_{I}$
    • Row: $x_{lj}=0\quad\forall j\in[1,N]$ if $j\neq m$ and $\forall(l,m)\in A_{I}$
    • Column: $x_{im}=0\quad\forall i\in[1,N]$ if $i\neq l$ and $\forall(l,m)\in A_{I}$
    • Region: $x_{ij}=0\quad\forall(i,j)\in R_{k}$ if $(i,j)\neq(l,m)$ and $(l,m)\in R_{k}$ $\forall(l,m)\in A_{I}$
In Fig. 2 a) we can see an example solution in the original N-queens, while in Fig. 2 b) we can see an example solution of the LinkedIn Queens.
Figure 2: a) Solution of an 8-queen problem, b) Solution of an 8-queen problem from LinkedIn
3 QUBO resolution
For the QUBO formulation of the problem, we need to be able to put the constraints as a quadratic form, making the constraints work for the variables in pairs. We first tackle the standard N-queens and then the LinkedIn queens.
For the N-queens we need all the $N^{2}$ variables, so we need $N^{2}$ logical qubits. The first constraint is an equality, which we can impose with a quadratic term of the form
$Q_{r}=\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}.$ (1)
This term has its minimum, equal to zero, when for each row there is only one variable equal to 1, so that all solutions that do not meet this restriction have a higher cost.
For the column condition, the term is analogous, interchanging the summands in $i$ and in $j$
$Q_{c}=\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (2)
For the restriction of the diagonals we use a term of the type
$Q_{d}=\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl},$ (3)
so that, for each occupied cell $(i,j)$, if another cell on its diagonal is occupied, this term sums to 1, otherwise it sums to 0. In this case we define $D^{\prime}_{i,j}$ as the cells that are on the same diagonals as cell $(i,j)$ such that their row is greater than $i$. In this way we avoid adding twice over the same pair of cells and we also avoid the term $x_{ij}^{2}=x_{ij}$, which would give us 1 whenever that cell was occupied. As only cells in rows below the cell where there is a queen are computed, we do not need to include the condition in the last row for $i$.
The total QUBO formulation is
$Q=$ $Q_{r}+Q_{c}+Q_{d}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}.$ (4)
All valid solutions to this problem have a cost $Q=0$, while all solutions that do not satisfy the constraints have a cost $Q>0$.
For the LQueens the formulation is similar. The first adjustment we have to make is to consider what are the fixed variables due to the initial condition. These are removed directly from the optimisation problem, but we keep the notation of the boxes for clarity. That is, if we have the $(3,3)$ box, the variable referring to the $(4,3)$ box will still be $x_{4,3}$, since we know that there really is no $x_{3,3}$. So we will have a number of variables equal to $N^{2}$ if there are no queens at the beginning, and less if we add queens in the initial condition.
Let’s define for simplicity the sets of cells:
  • $A_{C}$: contains the cells that have not been fixed by the initial condition (all variables to be optimised).
  • $A^{\prime}_{C}$: contains the cells of $A_{C}$ that are not in the last row of $A_{r}$.
  • $A_{R}$: contains the indexes of all but one of the regions that have at least one cell in $A_{C}$.
and the sets of indexes:
  • $A_{r}$: contains all rows $i$ containing at least one cell in $A_{C}$.
  • $A_{c}$: contains all columns $j$ containing at least one cell in $A_{C}$.
  • $A_{c}(i)$: the set of $j$ such that $(i,j)\in A_{C}$.
  • $A_{r}(j)$: the set of $i$ such that $(i,j)\in A_{C}$.
Thus, the term of the rows is
$Q_{r}=\sum_{i\in A_{r}}\left(1-\sum_{j\in A_{c}(i)}x_{ij}\right)^{2}.$ (5)
The term of the columns is
$Q_{c}=\sum_{j\in A_{c}}\left(1-\sum_{i\in A_{r}(j)}x_{ij}\right)^{2}.$ (6)
In both cases, the only modification with respect to the N-queens has been to eliminate from the summations the rows and columns without cells to be determined and the cells already fixed.
For the diagonal condition in this case, we have the same term, redefining in this case $D^{\prime}_{ij}$ as the diagonal cells adjacent to cell $(i,j)$, such that their row is greater than $i$, instead of the whole diagonals, and which have not already been fixed by the initial condition. Because of the initial condition, we have to construct the term as
$Q_{d}=\sum_{(i,j)\in A^{\prime}_{C}}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl},$ (7)
In this case we only consider cells that are not fixed.
Now, for the region condition, we add a term
$Q_{g}=\sum_{k\in A_{R}}\left(1-\sum_{(i,j)\in R_{k}|(i,j)\in A_{C}}x_{ij}\right)^{2}.$ (8)
ensuring that the minimum value 0 in the case that each region has only one queen in each region. We consider only $N-N_{I}-1$ regions, since if there is a queen in each of these, the remaining queen must necessarily be in the remaining region, and it can only be one, because of the above restrictions.
Therefore, the total QUBO formulation in this case is
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i\in A_{r}}\left(1-\sum_{j\in A_{c}(i)}x_{ij}\right)^{2}+$
$+$ $\sum_{j\in A_{c}}\left(1-\sum_{i\in A_{r}(j)}x_{ij}\right)^{2}+$
$+$ $\sum_{(i,j)\in A^{\prime}_{C}}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k\in A_{R}}\left(1-\sum_{(i,j)\in R_{k}|(i,j)\in A_{C}}x_{ij}\right)^{2}.$ (9)
As before, all valid solutions to this problem have a cost $Q=0$, while all solutions that do not satisfy the constraints have a cost $Q>0$.
4 General cases
In the previous section we have solved the particular cases of N-queens and the LQueens. However, we can generalise this formulation to solve more complex problems. As we have seen, each constraint has an associated QUBO term, and the initial condition determines the number of variables we have in the problem. Therefore, let’s determine the terms of the QUBO problem for each constraint. Since the constraints may be incompatible with each other, this formulation may also be used to test whether a certain generalized problem has a solution that satisfies its constraints.
As the initial condition only fixes the set of variables of the problem, in order to make the notation clear and without loss of generality, we will not add it. We will consider a rectangular board of dimensions $N\times M$. If we wanted irregular shapes, we would only have to create a rectangular board such that it can contain all the cells of the wanted one, and fix all cells outside the desired board to zero. In the same way, we can reduce a problem with an initial condition to a problem without an initial condition with a board with only the cells that are not fixed by the initialization, and adjusting the number of queens in each region according to the number of queens already existing in that region by the initial condition. We can see an example of an irregular board in Fig. 3 a), and how to translate a problem with initial condition to one without initial condition by changing the shape of the board in Fig. 3 b).
Figure 3: a) Irregular shaped board. b) Regular board with initial condition and its equivalent irregular board without initial condition. The cells in grey are squares that are not on the board.
4.1 Number of queens per row/column
The first generalised constraint to consider is the following. Given a board of dimension $N\times M$, there can only be $r_{i}$ queens in the $i$-th row. As can be understood, the row restriction in the cases studied above is the particular case where $r_{i}=1$ for all rows.
The term that implements this constraint is a generalisation of (1)
$Q_{r}=\sum_{i=1}^{N}\left(r_{i}-\sum_{j=1}^{M}x_{ij}\right)^{2}.$ (10)
The next constraint to implement is, given a board of dimension $N\times M$, there can only be $c_{j}$ queens in the $j$-th column. Once again, the column restriction in the cases studied above is the particular case where $c_{j}=1$ for all columns.
The term that implements this constraint is a generalisation of (2)
$Q_{c}=\sum_{j=1}^{M}\left(c_{j}-\sum_{i=1}^{N}x_{ij}\right)^{2}.$ (11)
In the case of an initial condition, these terms have to be modified both by eliminating the cells in which there can no longer be any queens (because their column and row already contain the required number of queens) and by reducing the number of queens required in the row and column in which we have already placed a queen. That is, if in row 3 we have placed a queen by initial condition, the number of queens available for the rest of cells in the row will be $r_{3}-1$, so in the QUBO we will have to reduce each $r_{i}$ and $c_{j}$ with the number of queens we initially placed in row $i$ and column $j$, respectively.
4.2 Distant diagonals constraint
In the case of N-queens, we required that if there was a queen on a cell, there could not be any more queens on any of its diagonals until the end of the board, whereas in LQueens it is only required for cells on the adjacent diagonal. To generalise this case, we say that we have a diagonal condition of distance $d_{ij}$ if in case there is a queen on the cell $(i,j)$ there cannot be any queen on any cell located on any of its diagonals at a distance equal to or less than $d_{ij}$ cells. In this way, not only do we allow diagonals of different lengths to be considered, but also that this depends on the cell.
The case of the N-queens is the particular case of $d_{ij}=\max(N,M)\quad\forall i\in[1,N],j\in[1,M]$, and LQueens that of $d_{ij}=1\quad\forall i\in[1,N],j\in[1,M]$.
The term that implements this is similar to the ones shown above in (3), as $d_{ij}$ only change the set $D_{i,j}$. In this case, because of the non-symmetry of the required diagonal distance between cells, $D^{\prime\prime}$ has to be used in the summation, instead of $D^{\prime}$. $D^{\prime\prime}_{ij}$ contains all the cells of $D_{ij}$ but $(i,j)$.
$Q_{d}=\sum_{i,j}^{N,M}\sum_{(k,l)\in D^{\prime\prime}_{i,j}}x_{ij}x_{kl},$ (12)
Given this abstract formulation, this condition can be generalized so that the 4 diagonals do not have the same distance, that instead of being continuous lines they can be skipped lines, that instead of being 4 diagonals they are $n_{d}$ lines with different slopes, or that they are directly abstract regions, since the only thing that change is the definition of the $D_{ij}$ over which to sum for each cell.
4.3 Number of queens per region
In the case of LQueens we imposed that, for a set of $N$ non-overlapping regions, in each of them there should be one and only one queen. However, we can generalize it so that regions can overlap, since it will only affect the definition of the $R_{k}$, and that instead of there having to be 1 queen per region, there must be $q_{k}$ queens in the $k$-th region.
This term will be a generalization of (8) for a number $N_{R}$ of regions.
$Q_{g}=\sum_{k=1}^{N_{R}}\left(q_{k}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (13)
As can be seen, this term is a generalization of both the row term and the column term, since we can make one region per row and one region per column, and it would result in the same terms.
If we had an initial condition, this term would have to be corrected by subtracting from $q_{k}$ the number of queens we have placed at the beginning in the $k$-th region, eliminating the cells that have already been fixed or that are in regions in which there can be no more queens.
Therefore, a generalized queens problem with a board of dimension $N\times M$ with no initial condition can be expressed by the two terms
$Q$ $=Q_{g}+Q_{d}=\sum_{k=1}^{N_{R}+N+M}\left(q_{k}-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N,M}\sum_{k,l\in D_{i,j}}x_{ij}x_{kl}$ (14)
where the set of regions $R^{\prime}$ is contains the $N_{R}$ regions $R_{k}$, the $N$ rows and the $M$ columns.
4.4 Two possible numbers of queens per region
A final generalization we can make is to create a constraint that is: given a board of dimension $N\times M$, in which there are $N_{V}$ regions that can be overlapping, there must be a number $v_{k}$ or $v_{k}+1$ of queens in region $k$ for each region. This constraint can be imposed by using a fractional term [8] that places the minimum cost in the middle of both values
$Q_{v}=\sum_{k=1}^{N_{V}}\left(v_{k}+\frac{1}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (15)
If we had an initial condition, we would have to apply the same correction as in the previous case.
As can be seen, the diagonal constraints are only particular cases of this constraint in which there is one region for each cell, which would be its diagonals. The value $v_{k}$ would be 0 for all the regions, since the number of queens must be 0 or 1. We can rewrite the eq (4.3), with $N_{R}$ regions $R_{k}$ with a single possible number of queens and $N_{V}$ regions $V_{k}$ with two consecutive possible numbers of queens, as
$Q=$ $Q_{g}+Q_{v}=\sum_{k=1}^{N_{R}}\left(q_{k}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}+$
$+$ $\sum_{k=1}^{N_{V}}\left(v_{k}+\frac{1}{2}-\sum_{(i,j)\in V_{k}}x_{ij}\right)^{2}.$ (16)
Finally, if we want to write the whole cost function as a single term, for $N_{R}$ regions $R_{k}$, we can define a value $t_{k}$, which will be 0 if the region $k$ can only have a single number of queens $q_{k}$ and 1 if the region $k$ can have two numbers of queens $q_{k}$ or $q_{k}+1$, such that
$Q=\sum_{k=1}^{N_{T}}\left(q_{k}+\frac{t_{k}}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (17)
5 Definition of generalised N-queens
With the terms and generalisations we have created for the N-queens, we can define a fully general N-queens, and its associated QUBO, and particularise it to a number of interesting families of N-queens problems.
The general N-queens problem is defined as:
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed $N_{I}$ initial queens in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which can overlap each other, such that in $k$-th region there must be $q_{k}$ queens if $t_{k}=0$ and $q_{k}$ or $q_{k}+1$ queens if $t_{k}=1$, and we have initially placed $p_{k}$ queens, the objective will be to place all the possible queens on the board such that the given restrictions are satisfied.
The QUBO formulation of this problem is given in general form as
$Q=\sum_{k=1}^{N_{R}}\left(q_{k}-p_{k}+\frac{t_{k}}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2},$ (18)
where $R_{k}$ is the set of cells included in the $k$-th region that have not been set in the initial condition and that do not belong to a region where the number of queens available after the initial queens have been placed is 0.
We will propose the form of two particular versions of this problem.
5.1 N-queens with regions
In this case the problem is a combination of the original N-queens with the LQueens, adding to the N-queens the $N_{R}$ regions in which there can only be one queen.
In this case, the QUBO formulation is
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{g}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{R}}\left(1-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (19)
5.2 N-queens with soft regions
This is the case of N-queens with regions, but allowing each region to have a queen or not.
In this case, the QUBO formulation is
$Q=$ $Q_{r}+Q_{c}+Q_{d}+Q_{v}=$
$=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{R}}\left(\frac{1}{2}-\sum_{(i,j)\in R_{k}}x_{ij}\right)^{2}.$ (20)
Another equivalent formulation would be to change the last term to one similar to that of the diagonals
$Q=$ $\sum_{i=1}^{N}\left(1-\sum_{j=1}^{N}x_{ij}\right)^{2}+\sum_{j=1}^{N}\left(1-\sum_{i=1}^{N}x_{ij}\right)^{2}+$
$+$ $\sum_{i,j}^{N-1,N}\sum_{(k,l)\in D^{\prime}_{i,j}}x_{ij}x_{kl}+$
$+$ $\sum_{k=1}^{N_{R}}\sum_{(i,j)\in R_{k}}\sum_{(l,m)\in R_{k}|(l,m)\neq(i,j)}x_{ij}x_{lm}.$ (21)
6 Chess Piece Problems
As can be seen, this formulation allows, not only to solve the case for the movement of queens, but also for the movement of any type of piece.
Therefore, we can generalise the problem to the chess piece problem, which will consist of trying to place as many pieces as possible on a board of maximum dimensions $N\times M$ so that no piece threatens another, with the extra restriction that there can only be one piece per coloured region and each cell can only be empty or have a specific type of piece, which can vary between cells.
The Coloured Chess Piece Problem is:
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, and divided into $N_{R}$ regions, which cannot overlap each other, such that in each region there must be only one piece, the objective will be to place all the possible pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}})_{ij}\right)$. A solution of this problem is represented in Fig. 4 a.
Figure 4: a) Solution of a rectangular Coloured Chess Piece Problem, b) Solution of a rectangular Max Chess Pieces Problem
As in the case of the queens, we will only have to eliminate the cells in which we have set the variables when initializing. That is, in which we have placed the pieces ($(i,j)\in A_{I}$), in the regions in which we have placed them ($(l,m)\in R_{k}|(i,j)\in R_{k},\forall(i,j)\in A_{I}$) and in the cells threatened by the placed pieces ($(l,m)\in C\left(M_{P_{ij}}\right)_{ij}\forall(i,j)\in A_{I}$). These cells can only be empty, so their variables will not be considered in the optimization, although we can place them in the QUBO formulation, since such values of 0 will not alter the summations. We will define the $R^{\prime}_{k}$ regions as the $N^{\prime}_{R}$ regions in which no pieces have been placed at the beginning.
This problem has the restrictions:
  • Coloured Regions: There can be only one piece per coloured region:
    $\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}=1\quad\forall k\in[0,N^{\prime}_{R}-1]$
  • No threat: No piece can be threatening another piece:
    $\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{lm}=0\quad\forall(i,j)|x_{ij}=1$
To implement these restrictions, we only need a coloured term
$Q_{g}=\sum_{k=0}^{N^{\prime}_{R}-1}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}$ (22)
and a threat term
$Q_{t}=\sum_{i,j=0}^{N-1,M-1}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (23)
The total QUBO formulation is
$Q=$ $Q_{g}+Q_{t}=\sum_{k=0}^{N^{\prime}_{R}-1}\left(1-\sum_{(i,j)\in R^{\prime}_{k}}x_{ij}\right)^{2}+$
$+\sum_{i,j=0}^{N-1,M-1}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (24)
Thus, a board that meets the constraints will be one with zero energy.
The other problem we can solve is the Max Chess Pieces Problem, which consists of placing as many pieces as possible on the board so that none of them threatens another, and there can be only one type per cell. We will define it as:
Given a board of arbitrary shape, of maximum height $N$ and maximum width $M$, on which there are placed a number $N_{p}$ of initial pieces in positions given by the set $A_{I}$, the objective will be to place the maximum number possible of pieces on the board such that no piece threatens another. Each cell $(i,j)$ can be empty or have the piece of type $P_{ij}$, with possible moves $M_{P_{ij}}$, and can reach cells $C\left(M_{P_{ij}})_{ij}\right)$. A solution of this problem is represented in Fig. 4 b.
Since we will only keep the threat constraint, we can eliminate the coloured term, and we have to include a new term that reduces the energy the more tiles there are on the board. This counting term is
$Q_{n}=-\sum_{i,j=1}^{N-1,M-1}x_{ij}$ (25)
so that the energy will decrease by one unit for each piece added to the board. This formulation even allows a different weight to be associated with each type of piece, so that it is more important to place some types of pieces than others.
To get the total QUBO, if we add directly the terms $Q_{t}$ and $Q_{m}$ we can have cases in which, by adding a single piece that threatens only one piece, we will have the same energy. Therefore, to make sure that the constraint is met, we will multiply the $Q_{t}$ term by a constant $\lambda$ large enough so that it is not convenient to add pieces that do not meet the constraint. Therefore, the total QUBO is
$Q=$ $Q_{n}+\lambda Q_{t}=-\sum_{i,j=1}^{N-1,M-1}x_{ij}+$
$+$ $\lambda\sum_{i,j=0}^{N-1,M-1}\sum_{(l,m)\in C\left(M_{P_{ij}}\right)_{ij}}x_{ij}x_{lm}.$ (26)
7 Conclusions
We have presented several generalizations of the N-queens problem, including the case of LinkedIn, presenting a general QUBO formulation to solve them, and interesting particular cases with their corresponding formulations. We have also presented two new interesting types of problems, associated with N-queens, and their respective QUBO formulations.
Future lines of research may include its implementation in various types of quantum hardware, benchmarking them using these problems, its extension to the case in which we allow more than two consecutive values for the number of queens in each region or its transformation to QUDO formulation.
Acknowledgment
This work is framed within the Quipucamayocs, the activities of the QuantumQuipu community.